OpenStudy (thefluffmuncher):

Solve the following equation by transforming it into a perfect square trinomial. x2 – 14x = 15

OpenStudy (thefluffmuncher):

How do i do this?

OpenStudy (thefluffmuncher):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

x^2 - 14 - 15 = 0 No need you can factor this out, (x )(x ) -15 * 1 = -15 -15 + 1 = -14 (x - 15)(x + 1) x = 15, -1

OpenStudy (thefluffmuncher):